OpenStudy (jordans_openstudy):

Write the point-slope form of the equation of a line with slope 2/5 that passes through the point (-3,-6).

OpenStudy (mathmale):

You have the slope: 2/5. You have the x-coordinate of the given point: -3 You have the y-coordinate also: -6 Substitute these values into the point-slope formula:\[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

Note: x_1 becomes -3. Leave the 'y' and 'x' as they are.

OpenStudy (jordans_openstudy):

ahh im kinda confused

OpenStudy (jordans_openstudy):

-24/5 is what my calculator shows @mathmale

OpenStudy (mathmale):

Your job is to come up with the equation of a straight line, and that equation must be in point-slope form. Your calculator won't give you this entire equation. I gave you the general point-slope equation:\[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

Here's how to get started: Take the given slope (2/5) and use this to replace m in the given equation. Type out your result here, please.

OpenStudy (jordans_openstudy):

y−y1=2/5 (x−x1) I haven't done this in a while. Where exactly do the -3,-6 go?

OpenStudy (mathmale):

In my very first response to your question, I typed You have the x-coordinate of the given point: -3 You have the y-coordinate also: -6 Please use this information. Finish this:\[y-y _{1}=(2/5)(x-x _{1})\]

OpenStudy (mathmale):

Leave x and y alone. Change only y_1 and x_1 by substituting the given values.

OpenStudy (mathmale):

Please, respond. Do something. Please get involved. If you knew this material once, you could learn it again with some effort.